OpenStudy (darkigloo):

is the series absolutely convergent, convergent, or divergent? (alternating series )

OpenStudy (darkigloo):

\[\sum_{n=1}^{\infty} -1^n n ^{-1}\ln(n+3) \] \[\sum_{n=1}^{\infty} \frac{ (n+2)^{n} }{ 2^{n}^{2} } \]

OpenStudy (darkigloo):

\[\sum_{n=1}^{\infty} \frac{ (n+2)^{n} }{ 2^{n^2}} \]

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \sum_{ n=1 }^{ \infty } (-1)^n\left[\color{blue}{\frac{\ln(n+3)}{n}}\right]}\)

OpenStudy (solomonzelman):

As I have said before, if there is a term n=j from which the series converge (like n=100, or n=98264 or any n=k), then this series will as well converge starting from n=1. (provided that none of the terms are singular)

OpenStudy (solomonzelman):

If you start from n=1000 for example, then you have a monotonically decreasing sequence \(A_n\), and the series alternates.

OpenStudy (solomonzelman):

So starting from n=1000 the series would converge, and therefore .....

OpenStudy (solomonzelman):

But, the question is, would this series converge without alternation?

OpenStudy (solomonzelman):

if it does without alternation, then it converges absolutely. if it does NOT converge without alternation, then it converges conditionally.

OpenStudy (solomonzelman):

Will \(\color{#000000}{ \displaystyle \sum_{ n=1 }^{ \infty } \left[\color{blue}{\frac{\ln(n+3)}{n}}\right]}\) converge ?

OpenStudy (solomonzelman):

As n tends to infinity you know that: \(\color{#000000}{ \displaystyle \color{blue}{\frac{(n+2)^n}{2^{n^2}}}>\color{blue}{\frac{(n+n)^n}{2^{n^2}}} }\) \(\color{#000000}{ \displaystyle \color{blue}{\frac{(n+n)^n}{2^{n^2}}} =\color{blue}{\frac{(2n)^n}{2^{n^2}}} =\color{blue}{\frac{2^nn^n}{2^{n^2}}}=\color{blue}{\frac{n^n}{2^{n^2}2^{-n}}}=\color{blue}{\frac{n^n}{2^{(n-1)n}}} }\) \(\color{#000000}{ \displaystyle =\color{blue}{\frac{n^n}{\left(2^{n-1}\right)^n}} }\)

OpenStudy (solomonzelman):

Oh my sign is wrong

OpenStudy (solomonzelman):

the first equation should say <. a simple typo, sorry

OpenStudy (solomonzelman):

So.... Would you agree that \(\color{#000000}{ \displaystyle \sum_{n=j}^\infty \left(\frac{n}{2^{n-1}}\right)=\rm converges }\) ? Now, since \(\color{#000000}{ \displaystyle \left(\frac{n}{2^{n-1}}\right)<1 }\) therefore \(\color{#000000}{ \displaystyle \left(\frac{n}{2^{n-1}}\right)>\left(\frac{n}{2^{n-1}}\right)^n }\) (Because raising number x to larger powers if |x|<1 makes the result smaller) and so, all the more so \(\color{#000000}{ \displaystyle \sum_{n=j}^\infty \left(\frac{n}{2^{n-1}}\right)^n=\rm converges }\) (because this series is smaller than the previous) and since (as I deduced) \(\color{#000000}{ \displaystyle \frac{(n+2)^n}{2^{n^2}}>\left(\frac{n}{2^{n-1}}\right)^n }\) therefore Would you agree that \(\color{#000000}{ \displaystyle \sum_{n=j}^\infty \left(\frac{n}{2^{n-1}}\right)=\rm converges }\) Since \(\color{#000000}{ \displaystyle \left(\frac{n}{2^{n-1}}\right)<1 }\) therefore \(\color{#000000}{ \displaystyle \left(\frac{n}{2^{n-1}}\right)>\left(\frac{n}{2^{n-1}}\right)^n }\) and so, all the more so \(\color{#000000}{ \displaystyle \sum_{n=j}^\infty \left(\frac{n}{2^{n-1}}\right)^n=\rm converges }\) and since (as I deduced) \(\color{#000000}{ \displaystyle \frac{(n+2)^n}{2^{n^2}}>\left(\frac{n}{2^{n-1}}\right)^n }\) therefore \(\color{#000000}{ \displaystyle \sum_{n=j}^\infty \frac{(n+2)^n}{2^{n^2}}=\rm converges }\) ! ■

OpenStudy (solomonzelman):

and the first series converges, like any polynomial over exponential. In particular, you can start from n=1000000, and compare the polynomial numerator to 1.5^n (because at that starting point,and on forever, 1.5^n is greater than your polynomial), and thus you get (1.5^n/2^n)=(1.5/2)^n so your series is behaving smaller than a geometric series with |r|<1.

OpenStudy (solomonzelman):

I said something twice there.... hope it's not too confusing. good luck.

satellite73 (satellite73):

i still can't figure out what the series is

satellite73 (satellite73):

is it this: \[\sum_{n=1}^{\infty} \frac{ (n+2)^{n} }{ 2^{n^2}}\]?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

and the first one as well.

OpenStudy (zarkon):

\(\displaystyle \sqrt[n]{\frac{ (n+2)^{n} }{ 2^{n^2}}}=\frac{n+2}{2^n}\to 0<1\) as \(n\to\infty\)

OpenStudy (solomonzelman):

oh, I was overthought that:) N-th root test:)